### v2

# Free-Lunch Color-Texture Disentanglement for Stylized Image Generation

Abstract
Recent advances in Text-to-Image (T2I) diffusion models have transformed image generation, enabling significant progress in stylized generation using only a few style reference images. However, current diffusion-based methods struggle with fine-grained style customization due to challenges in controlling multiple style attributes, such as color and texture. This paper introduces the first tuning-free approach to achieve free-lunch color-texture disentanglement in stylized T2I generation, addressing the need for independently controlled style elements for the Disentangled Stylized Image Generation (DisIG) problem. Our approach leverages the Image-Prompt Additivity property in the CLIP image embedding space to develop techniques for separating and extracting Color-Texture Embeddings (CTE) from individual color and texture reference images. To ensure that the color palette of the generated image aligns closely with the color reference, we apply a whitening and coloring transformation to enhance color consistency. Additionally, to prevent texture loss due to the signal-leak bias inherent in diffusion training, we introduce a noise term that preserves textural fidelity during the Regularized Whitening and Coloring Transformation (RegWCT). Through these methods, our Style Attributes Disentanglement approach (SADis) delivers a more precise and customizable solution for stylized image generation. Experiments on images from the WikiArt and StyleDrop datasets demonstrate that, both qualitatively and quantitatively, SADis surpasses state-of-the-art stylization methods in the DisIG task. Code will be released at here.
{strip}[Uncaptioned image]
Figure 1: Stylized images generated by our training-free method, SADis. (Up) As shown in the first two rows, it enables disentangled control over color and style attributes in text-to-image diffusion models using separate image prompts. This approach offers creators enhanced color control, including the use of color palettes as in the last two columns. (Down) SADis also enables real-image stylization by incorporating a content image as an additional condition via ControlNet. Furthermore, it extends to color-only stylized generation and material transfer for more flexible image generation.
1 Introduction
Stylized Image Generation [12, 30, 36], also termed style transfer in classical research, aims to transfer a style from a reference image to a target image. This field has evolved through several technological paradigms, beginning with CNN-based feature manipulation [20, 65, 36], advancing to attention mechanisms [4, 74, 46], and further developing through GAN-based image translation [37, 81, 72]. A significant advancement came with multimodal CLIP guidance [18, 40, 52], which bridged the gap between visual and textual representations. This breakthrough enabled a novel approach where style transfer could be guided by textual descriptions rather than being limited to reference images [18, 41, 35]. The field underwent another transformation with the emergence of Text-to-Image (T2I) diffusion models [56, 59, 57], which demonstrated remarkable capabilities in personalized image generation [58, 19, 49, 7].
T2I diffusion models catalyzed new developments in stylized image generation [54, 67, 62] — a specialized paradigm that focuses on creating new images that incorporate specific visual characteristics from reference styles, rather than simply transferring style between existing images. Initially, many approaches [19, 2] relied on extensive fine-tuning using datasets of similarly styled images. However, this requirement proved impractical in real-world scenarios, where collecting cohesive style-specific datasets is often challenging. Addressing these limitations, recent research has focused on developing tuning-free (free-lunch) methods [75, 54, 38]. These approaches eliminate the need for costly retraining while maintaining efficient style integration capabilities. Despite their methodological differences, both tuning-based and tuning-free approaches share a common characteristic: they transfer entangled style representations during the image generation process, simultaneously incorporating both color and texture elements from the style images into the final output.
Although existing stylized image generation methods offer promising advancements, for content creators, these methods have critical limitations. A key challenge is the lack of granular control over style elements. Content creators often need to adopt specific aspects of a reference style while preserving elements of their original vision [14, 27]. Color palettes, in particular, play a crucial role in this process — they are often meticulously crafted to evoke specific emotions or maintain brand consistency. Creators may wish to preserve these while adopting textural [^1] In this paper, we use the term texture to refer to the arrangement, repetition, and local patterns of visual elements such as shapes, edges, intensities, and gradients and excluding color aspects. elements from other reference styles. However, current approaches force an all-or-nothing choice, where accepting a reference style means incorporating all its visual (texture and color) characteristics simultaneously, seriously limiting artistic freedom and practical applicability.
To address these limitations, we introduce the problem of Disentangled Stylized Image Generation (DisIG), which aims to decompose reference styles into independently controllable attributes like color, texture, and semantic content. This formulation enables artists to selectively transfer specific style elements while maintaining control over the content using text prompts. One potential workaround for DisIG using existing methods is to specify desired style attributes through text prompts. However, this approach requires extensive prompt engineering [48, 76, 71] expertise and considerable trial-and-error to achieve results that can be obtained more intuitively through image prompts. Even with significant effort in crafting precise textual descriptions, text prompts often fall short of capturing the nuanced characteristics of the target style. This limitation stems from the inherent expressiveness gap between text and visual information — image prompts inherently contain richer semantic information, including style attributes such as color and texture, than text prompt descriptions, that nuanced details lead to enhance generation quality as shown in Fig. 2.
In this work, we propose what we believe to be the first DisIG method that enables independent control of color and texture through separate reference images, without any training. To achieve this, we first analyze the additivity property — a characteristic previously studied in textual spaces [51, 6, 33] — and demonstrate that it also holds true in the image prompt space. Then, we work with each style component separately: we isolate the color representation through feature subtraction between the original color image and its grayscale equivalent, while deriving the texture representation using a grayscale version of the texture reference image and SVD rescaling. Finally, to ensure accurate color matching while preserving textural details, we introduce RegWCT, an enhanced Whitening and Coloring Transformation with noise regularization that counteracts the signal-leak bias [16, 80] common in diffusion models.
In experiments, we use images from the WikiArt [64] and StyleDrop [62] datasets and evaluate our approach based on the SDXL model [53]. By comparing against several state-of-the-art stylization methods, SADis consistently outperforms these baselines over both qualitative and quantitative results, particularly in achieving accurate color and texture expression for the stylized T2I generation. In summary, our contributions are as follows:
  • We introduce the Disentangled Stylized Image Generation (DisIG) problem, a critical challenge in real-world applications that enables richer semantic control in diffusion-based stylization.
  • We are the first to identify and leverage the image-prompt additivity property in the CLIP image encoder, showing how image representations can be decomposed and combined similarly to text embeddings.
  • We present the first tuning-free (free-lunch) style attribute disentanglement method, SADis, emphasizing color and texture as key style attributes. Our color-texture extraction (CTE) and regularized whitening-coloring transformation (RegWCT) techniques enable effective disentangled color-texture stylization without additional tuning.
  • Through comprehensive qualitative and quantitative evaluations in standard style transfer benchmarks, SADis consistently outperforms state-of-the-art baselines, especially in realistic T2I generation under the DisIG scenario.
Figure 2: Compared to existing stylization methods (InstantStyle and DEADiff), our SADis offers flexibility in style control by enabling the use of separate color and texture reference images [^3] Here we use the prompt “A ballerina” as an example. The prompt used in InstantStyle and DEADiff is a detailed description of the color reference image generated by GPT-4o [1]..
2 Related Work
Stylized image generation [30, 12, 37, 81], also related to classical style transfer, aims to generate images with the artistic style of a reference image. Early research focused on statistical feature manipulation to transfer artistic styles from reference images. Gatys et al. [20] pioneered this direction by using covariance matrices for style representation. Following works like AdaIN [36] improved efficiency by transferring feature statistics between style and content feature maps, while WCT [44] introduced whitening and coloring transformations to match covariance matrices. Recent architectural innovations have enhanced stylization capabilities. A significant paradigm shift occurred with the introduction of CLIP [55], enabling connections between text and image representations in a shared embedding space. This advancement spawned methods like ClipStyler [41], which combines global and patch-level CLIP losses for high-resolution stylization.
Figure 3: Our method, SADis, begins with color-texture extraction (CTE), leveraging the Image-Prompt Additivity property, which is verified in this paper for the first time. After obtaining the disentangled color and texture embeddings from the reference images, we incorporate these embeddings into the style cross-attention layer of the SDXL model. Subsequently, we refine the latent variable $z_{t-1}$ at each inference step with our proposed RegWCT transformation, aligning color palettes precisely while retaining essential texture details.
Nowadays, T2I Diffusion models [61, 29, 9] have emerged as the new state-of-the-art for text-to-image generation [8, 45, 33, 47]. Then the existing T2I stylization methods [23, 60, 13, 78, 66, 22, 79] achieve stylized image generation via fine-tuning generative models on few style reference images. However, this process is time-consuming and struggles to generalize to real-world scenarios where gathering a suitable subset of shared-style images can be difficult. To address these limitations, interest in tuning-free (free-lunch) approaches for stylized image generation has grown [11, 15, 70, 69, 26, 24]. These methods introduce lightweight adapters that extract style information from reference images and inject it into the diffusion process via self-attention or cross-attention layers. Representative examples include IP-Adapter [75] and Style-Adapter [69], which employ a decoupled cross-attention mechanism that separates cross-attention layers for handling text and image features independently. DEADiff [54] introduces a different approach by focusing on extracting disentangled representations of content and style using paired datasets.
Although existing stylized image generation methods show promise, they lack granular control over individual style elements - a crucial need for content creators who often want to selectively adopt specific aspects of reference styles. This introduces a challenge we term the Disentangled Stylized Image Generation (DisIG). Focusing on color and texture as the most significant style attributes, this paper introduces the first free-lunch approach to color-texture disentanglement for stylized image generation.
3 Method
3.1 Preliminaries
Problem Setup. Unlike classical stylized image generation [44, 10, 40], which derives style information from a single reference image, Disentangled Stylized Image Generation (DisIG) leverages two distinct reference images: a color reference image $\mathcal{I}_{clr}$, and a texture reference image $\mathcal{I}_{tx}$. The color reference image could either be a normal RGB image or a color palette. Given an input prompt $\mathcal{P}$ for the T2I model, our goal is to generate an image aligning with the prompt while exhibiting a color palette similar to $\mathcal{I}_{clr}$ and texture from $\mathcal{I}_{tx}$.
T2I Diffusion Models. We built on the SDXL [53] model, consisting of two primary components: an autoencoder and a diffusion model $\epsilon_{\theta}(z_{t},t,\tau_{\xi}(\mathcal{P}))$, where $\epsilon_{\theta}$ is a UNet, conditioning a latent input $z_{t}$, a timestep $t\sim\text{U}(1,T)$, and a text embedding $\tau_{\xi}(\mathcal{P})$. More specifically, text-guided diffusion models generate an image from the textual condition as $\mathcal{C}_{text}=\tau_{\xi}(\mathcal{P})$, where $\tau_{\xi}$ is the CLIP text encoder [55] [^4] SDXL uses two text encoders and concatenate the embeddings.. The cross-attention map is derived from $\epsilon_{\theta}(z_{t},t,\mathcal{C}_{text})$. After predicting the noise, diffusion schedulers [63, 50] are used to predict the latent $z_{t-1}$. Here we simplify the $z_{t-1}$ inference process as $z_{t-1}=\mathcal{G}(z_{t},t,\mathcal{C}_{text})$.
IP-Adapter. Building on T2I diffusion models, the IP-Adapter [75] introduces additional controllability by conditioning the T2I model on a conditional image $\mathcal{I}_{ip}$. Practically, this involves leveraging a pre-trained T2I diffusion model and incorporating a cross-attention layer to the (projected) image condition following each text-prompt conditioning layer. The conditional image is encoded in the low-dimensional CLIP image embedding space [55] to capture high-level semantic information. By denoting the CLIP image encoder as $\tau_{\phi}$ and IP-Adapter projection as $\mathbf{IP}$, this process is adding a new image condition $\mathcal{C}_{img}=\mathbf{IP}(\tau_{\phi}(\mathcal{I}_{ip}))$ to the T2I model as $z_{t-1}=\mathcal{G}(z_{t},t,\mathcal{C}_{text},\mathcal{C}_{img})$.
Figure 4: By subtracting the grayscale image embeddings from the color image embeddings, we can effectively remove the texture information and isolate the color information. Combined with the prompt “A photo of a bird” we can generate images in the same color schemes as the input images.
3.2 Style Attributes Disentanglement
Here, we present our Style Attributes Disentanglement method (SADis). An overview is provided in Fig. 3. To successfully extract the color and texture information, we exploit the image-prompt additivity property, explained in Sec. 3.2.1. Next, we explain the color-texture extraction (CTE) in Sec. 3.2.2, and the noise-regularized whitening-coloring transformation (RegWCT) is detailed in Sec. 3.2.3.
3.2.1 Image-Prompt Additivity
To achieve disentangled stylization, we first analyze the additivity property within image prompts, which can be derived using the CLIP image encoder, as also adopted in the IP-Adapter [75]. This property, which has been explored in text prompt spaces [51, 6], also holds in the image prompt space. As an example in Fig. 4-(Left), we subtract the grayscale image embedding from the original color image embedding, that results in the T2I generation with similar color palettes. We show additional analysis with subtracting objects from the given images in the Supplementary Material. We refer to this property as Image-Prompt Additivity. Using this property, we extract color and texture information from reference images and apply these attributes effectively in T2I generation.
3.2.2 Color-Texture Extraction
Leveraging this image-prompt additivity, we disentangle color and texture representations from their respective reference images. For color extraction, we take the color image embedding and subtract the grayscale image embedding:
$\mathbf{Emb}_{clr}=\tau_{\phi}(\mathcal{I}_{clr})\;\text{--⃝}\;\tau_{\phi}(\mathbf{GS}(\mathcal{I}_{clr}))$ (1)
where $\mathbf{GS}$ is the grayscale operation. This strips away semantic information to retain only the color attributes.
For texture extraction, we first convert the texture reference image to grayscale to remove any influence of its color palette, which is formulated as $\mathbf{Emb}_{tx}^{*}=\tau_{\phi}(\mathbf{GS}(\mathcal{I}_{tx}))$, However, using only the grayscale texture embedding creates a mismatch in scale between the color and texture branches, often resulting in overly gray tones in T2I generations (as shown in Fig. 5-2nd col.). To address this, we average the gray texture image to obtain a pure gray image, which summarizes the main gray tone information in the $\mathbf{GS}(\mathcal{I}_{tx})$ image. Following that, we concatenate both embeddings to obtain the initial texture representations:
[…]
Next, we conduct the Singular-Value Decomposition (SVD) over the texture representations. Inspired by [21, 43], we assume that the main singular values of $\mathbf{Emb}_{tx}^{{}^{\prime}}$ correspond to the shared fundamental information of these two grayscale images, specifically the grayscale tone. We then have: $\mathbf{Emb}_{tx}^{{}^{\prime}}=\textbf{\emph{U}}{\boldsymbol{\Sigma}}{\textbf{\emph{V}}}^{T}$, where $\boldsymbol{\Sigma}=diag(\sigma_{0},\sigma_{1},\cdots,\sigma_{n_{j}})$, the singular values $\boldsymbol{\sigma}_{0}\geq\cdots\geq\boldsymbol{\sigma}_{n_{j}}$ To suppress the gray tone expression and extract the texture information, we introduce the augmentation for each singular value as:
$\hat{\sigma}={\beta e}^{-\gamma\sigma}*\sigma.$ (3)
where $e$ is the exponential, $\gamma$ and $\beta$ are parameters with positive numbers. We then recover it as $\mathbf{Emb}_{tx}=\textbf{\emph{U}}{\hat{\boldsymbol{\Sigma}}}{\textbf{\emph{V}}}^{T}$, with the updated $\hat{\boldsymbol{\Sigma}}=diag(\hat{\sigma_{0}},\hat{\sigma_{1}},\cdots,\hat{\sigma_{n_{j}}})$. That effectively reduces residual gray-color influences while preserving texture details. After extracting both color and texture reference image embeddings, we achieve a baseline of the DisIG problem. The T2I inference process is now formulated as: $z_{t-1}^{{}^{\prime}}=\mathcal{G}(z_{t},t,\mathcal{C}_{text},\mathbf{Emb}_{tx}\oplus\mathbf{Emb}_{clr})$. Note that, we only inject the disentangled embedding $\mathbf{Emb}_{tx}\oplus\mathbf{Emb}_{clr}$ into the first decoder layer to compute cross-attention maps, which shows better stylization performance as previous works proved [67, 68, 2, 66]. Examples shown in Fig. 5 (2nd and 3rd cols.) demonstrate the effectiveness of our proposal.
Figure 5: We suppress the gray expression by introducing extra SVD reweighting to the texture representations.
3.2.3 Regularized Whitening-Coloring Transforms
WCT: whitening-coloring transformation. Solely applying our CTE method does not fully ensure a precise color palette match between the reference color image $\mathcal{I}_{clr}$ and the generated output. We hypothesize that while CLIP embeddings capture high-level semantic information, they may not fully represent nuanced color distribution details. To better align the color distributions of the generated image with the color reference, we incorporate a whitening-coloring transform (WCT) [44, 10, 31] applied to the noisy latent as $z_{t}^{wct}=\mathbf{WCT}(z_{t}^{{}^{\prime}})$. Please refer to the Supplementary for detailed WCT formulas. While traditionally used in style transfer, we are the first to apply this transformation directly within the latent space of a T2I diffusion model. This approach allows for a more faithful color transfer by aligning the latent representations with the reference color palette, enhancing the stylistic consistency of generated images. Examples of applying the WCT transform during the T2I generation process are presented in Fig. 6-(a), demonstrating an improved alignment of color palettes between the generated images and the reference color images. This enhancement visually confirms the effectiveness of WCT in achieving more accurate color transfer, resulting in a closer match to the desired color distribution.
Noise Regularization for WCT transform. However, as shown in Fig. 6-(a), the WCT process tends to compress texture details, distorting high-frequency information, which is also observed in previous works [80, 42]. This issue is attributed to the signal-leak bias [16] and the inverted latent distribution gap from the inversion process [80]. These terms essentially describe the same underlying phenomenon: during training, the distribution of $z_{T}$—the latent representation after adding Gaussian noise to $z_{0}$—does not perfectly match a standard Gaussian. Instead, it always contains the leakage or some certain prior information towards the original image, causing the residual structure to persist even at high noise levels. This biased noise in the forward diffusion process leads to a slight misalignment in texture representation during generation. To address this, we propose adding a small-scale noise to the latent representation as $z_{t}=z_{t}^{wct}+\lambda\cdot\delta,\delta\sim\mathcal{N}(0,1)$, determined by a scale hyperparameter $\lambda$, to recover the lost high-frequency details. By introducing this noise regularization term, the regularized whitening-coloring transformation (RegWCT) achieves improved generative performance, better aligning both color and texture in the generated images with the reference images. Given a latent $z_{t}^{{}^{\prime}}$ at timestep $t$, the color rectified latent $z_{t}$ after RegWCT is formulated as $z_{t}=\left(1-\omega\right)z_{t}^{{}^{\prime}}+\omega\cdot\mathbf{RegWCT}(z_{t}^{{}^{\prime}})$, where $\omega$ acts as a balance weight. To prevent the inference trajectory from deviating too much from the original one, we only perform the RegWCT transformation during the intermediate steps as $[T_{start},T_{end}]$. With all these techniques, including color-texture extraction (CTE) and Regularized WCT transformation (RegWCT), our method SADis achieves customizable and flexible Disentangled Stylized Image Generation (as shown in Fig. 5-4th col.).
Figure 6: (top) We enable preciser color alignment with $\mathcal{I}_{clr}$ through RegWCT, while WCT manipulation compresses texture details. (bottom) The distribution of the image latent generated with RegWCT technique is closer to that of the image generated without WCT in the frequency domain, across diverse time steps.
3.2.4 ControlNet-based real image stylization
Furthermore, our method SADis can be integrated with ControlNet [77] $\mathcal{CN}$ to facilitate image-based stylized generation, as illustrated in Fig. 11 (Left). By utilizing any pretrained ControlNet model (e.g. Canny-conditioned) as the base pipeline while maintaining all other hyperparameters consistent with SADis, we are able to significantly broaden the applicability of our method SADis to real image cases. More specifically, we have an input $\mathbf{I}_{c}$ as the content image, it is passed through the ControlNet $\mathcal{CN}(\mathbf{I}_{c})$ as conditions for T2I generation model as $z_{t-1}^{{}^{\prime}}=\mathcal{G}(z_{t},t,\mathcal{CN}(\mathbf{I}_{c}),\mathbf{Emb}_{tx}\oplus\mathbf{Emb}_{clr})$, where $z_{T}\sim\mathcal{N}(0,1)$ and the textual prompt as null $\mathcal{P}=\text{``''}$ “”.
Figure 7: Disentangled Stylized Image Generation (DisIG) performance compared with baseline methods. For other approaches, we used GPT-4o to generate color descriptions based on the color reference image, incorporating these descriptions into the text prompts (details shown in the Supplementary). Our SADis accepts separate color and texture reference images to achieve flexible control.
4 Experiments
4.1 Experimental Setups
Datasets. To ensure a fair comparison, we randomly select 40 images in total from the WikiArt [64] and StyleDrop [62] datasets. For our method, SADis, each of these images serves as either a color reference or a texture reference to enable color-texture disentanglement. In contrast, since the comparison methods lack disentanglement capabilities, we treat each of these 40 images as a style reference for these methods, supplemented by auxiliary text prompts generated from GPT-4o [1] as a strong captioning model. For quantitative comparisons, we used 20 images as the color reference set and 20 as the texture reference set. We also sampled 10 content prompts from StyleDrop, resulting in 4,000 stylized images per method to ensure fair and extensive comparisons with numerous images.
Evaluation Metrics. We evaluate our method using multiple quantitative metrics to ensure a comprehensive assessment. (1) For a general quality evaluation, we use CLIP-Score (CLIP) [28] to evaluate the T2I generation performance, specifically measuring the semantic consistency between the generated image and the text prompt. (2) For the evaluation of color attribute alignment, the MS-SWD [25] metric specifically evaluates color distance between generated and reference images. Color histogram distance (C-Hist) computes the distance between both color histograms. The GPT-4o color score [1] requests the GPT model to rate from 0-5 according to the color consistency between the color reference and the generated image. (3) For the texture quality, the Kernel-Inception Distance (KID) [5] is assessing the quality of generated images by measuring the dissimilarity between the real and generated image distributions [^5] We use feature dimension as 2048 and evaluate on the 4000 images.. CLIP-I is used to evaluate the similarity between the texture image and the generated image in grayscale.
Figure 8: SADis remains effective when exchanging texture and color images (left) or varying color image properties such as saturation and illuminant (right). This further demonstrates the robustness of SADis in realistic DisIG scenario.
Implementation Details. We build our method SADis upon the SDXL [53]. To apply the image embeddings as additional conditions to the T2I model, we use the IP-Adapter [75] pretrained projectors. Both of these models are based on the CLIP model [55], where the SDXL model utilizes the CLIP text encoder to generate textual embeddings and the IP-Adapter leverages the CLIP image encoder to extract image embeddings before the embedding projector. For the hyperparameters, we set $\lambda=0.01,T_{start}=0.8T,T_{end}=0.6T,\omega=0.5,\gamma=0.003,\beta=1.0$.
Comparison Methods. We compare with several state-of-the-art tuning-free stylization approaches based on the T2I diffusion models: DEADiff [54], InstantStyle [67], IP-Adapter [75], CSGO [73], DreamStyler [3], StyleDrop [62] and Artist [39]. For these baseline methods, we utilize the GPT-4o [1] model to generate captions from the color reference image, which are appended to the original textual prompt to guide the T2I generation. Additionally, SDXL [53] and Artist model are included in comparisons by generating captions from both the color and texture reference images, appending these captions to the textual prompts to guide the generation process. We include much more experimental setup details, qualitative and quantitative results in the Supplmentary Material.
4.2 Experimental Results
Qualitative Comparison. In our comparison of generation quality, visual results under the Disentangled Stylized Image Generation (DisIG) scenario are shown in Fig. 7 and  Fig. 9. While baseline methods manage to capture some style information from the style reference image (the texture image for SADis), they fail to represent the color palette provided by the color reference accurately. The IP-Adapter and InstantStyle methods even retain original layouts and elements from the style images, deviating from the intended stylization. Other baselines, such as Artist [39] and DEADiff [54], exhibit poor generation quality, reflecting a limited understanding of both color and texture attributes. In contrast, our method, SADis, allows for both texture and color reference inputs, offering precise, controllable T2I stylization. The generated images by SADis exhibit significantly improved color fidelity and texture detail compared to other approaches, demonstrating its effectiveness in disentangled style attribute transfer. To further validate SADis’s robustness, we conduct two tests: (1) swapping the color and texture reference images, and (2) modifying their saturation and illumination levels, as shown in Fig. 8. Unlike existing approaches such as InstantStyle, which cannot preserve image content under these conditions (detailed in Supplementary), SADis successfully applies color and texture independently while keeping the content largely unchanged.
Method CLIP Color Texture
[…]
SDXL [53] 0.272 9.51 1.20 3.01 0.69 0.08
IPAdapter [75] 0.233 11.54 1.23 2.84 0.84 0.043
InstantStyle [67] 0.261 12.53 1.32 2.80 0.74 0.056
Artist [39] 0.269 10.48 1.39 2.82 0.69 0.089
DEADiff [54] 0.267 11.20 1.24 2.73 0.69 0.087
StyleDrop [62] 0.275 13.52 1.43 2.67 0.70 0.054
DreamStyler [3] 0.277 12.17 1.26 2.39 0.71 0.060
CSGO [73] 0.280 14.25 1.36 2.63 0.69 0.071
SADis (Ours) 0.281 5.57 0.96 3.34 0.74 0.049
Table 1: Quantitative Comparison. The best and second-best numbers are marked with bold and underlined respectively.
Method Color Texture
[…]
SADis (Ours) 5.57 0.96 0.74
$\boldsymbol{-}$ SVD 5.70 1.01 0.76
$\boldsymbol{-}$ RegWCT 8.05 1.06 0.75
$\boldsymbol{-}$ SVD $\boldsymbol{-}$ RegWCT 8.93 1.10 0.76
Table 2: Ablation by removing each components of SADis.
Figure 9: SADis performs well using the same image as both the color and texture reference, demonstrating its flexible capability.
Figure 10: User study on comparison methods: evaluating color, texture, and combined color&textue alignment.
Quantitative Comparison. A detailed quantitative comparison is provided in Tab. 1 to further support our findings. Our method, SADis, retains text-image alignment quality at a level comparable to the base SDXL model [53], as indicated by the CLIP score. This comparison demonstrates that SADis preserves alignment with the textual prompt better than other methods. Regarding color alignment, SADis significantly outperforms other approaches. For texture representation, the IP-Adapter-based methods, including InstantStyle [67] and SADis, show superior performance over other baselines. They achieve higher texture-related metrics at the expense of color fidelity and text-image alignment quality. The high texture scores for IP-Adapter can be explained by the fact that they control all cross-attention layers; However, this leads to high-semantic content leakage (like the same stars in the Van Gogh bear in Fig. 7). We decided to prevent this by only controlling the cross-attention of the low-level layers following InstantStyle [67, 2, 66]; Therefore, our texture results are close to those of InstantStyle, but at the same time we greatly improve color fidelity. In addition, their generated images follow the same structure as the original texture image, showing that they do not disentangle structure from texture but overfit to the contents of the texture images.
Ablation Study. The ablation study for each component of SADis is listed in Tab. 2. Results show that SVD rescaling in the CTE process and the noise-regularized RegWCT techniques significantly enhance color alignment while only slightly reducing texture precision, which is almost undetectable in the CLIP-I metric. The trade-off between texture and color alignment is optional for users, allowing them to adjust the balance based on the requirements.
User Study. To better assess alignment with human preferences, as shown in Fig. 10, we conducted a user study with 24 participants (30 sextuplets/user), collecting 720 data for each method. Each one was asked to “select the best image from pairs of generated images, taking into account overall quality (considering both color and texture alignments), color alignment, and texture alignment, respectively.” Our method, SADis, significantly outperformed other baseline approaches, with at least a 20% improvement in preference. That finally demonstrates the potency of SADis and high alignment with human preference.
4.3 Additional Applications
Image-based stylization. SADis Integrated with ControlNet achieves image-based stylized generation, as shown in Fig. 11(left). Here, we adopt ControlNet (Canny) as the base model, setting its conditioning scale to 0.6. This integration broadens the application scenarios of SADis.
Figure 11: SADis allows for real-image based stylization (left). Besides, SADis effectively disentangles color and material elements from separate images, enabling precise control over color and material in image generation.
Color and material transfer. Given color and material references from separate images, SADis is also capable of disentangling color and material elements, and provides precise controls over color and materials in T2I generation. The color and material transfer results are shown in Fig. 11 (right). This capability of SADis highlights its potential applications in artistic creation and industrial design.
5 Conclusion
In this paper, we address a critical challenge in stylized image generation by introducing the concept of disentangled stylized image generation (DisIG). We focus on color and texture as key style attributes, presenting the first approach, named style attribute disentanglement (SADis), for independent control over these elements. Using the image-prompt additivity property, we proposed novel techniques, including the color-texture extraction (CTE) and regularized whitening-coloring transformation (RegWCT), to ensure enhanced color-texture consistency and more accurate results. Experimental evaluations demonstrate that SADis significantly outperforms existing stylization methods, both qualitatively and quantitatively. This work not only advances the field of T2I stylization but also opens new avenues for more flexible and customizable image generation, paving the way for future innovations for art creators.
Acknowledgements
We acknowledge project PID2022-143257NB-I00, financed by MCIN/AEI/10.13039/501100011033 and FEDER, and the Generalitat de Catalunya CERCA Program. This work was also supported by NSFC (NO. 62225604) and Youth Foundation (62202243). We acknowledge “Science and Technology Yongjiang 2035” key technology breakthrough plan project (2024Z120).
Supplementary Material: Free-Lunch Color-Texture Disentanglement for Stylized Image Generation
Appendix S1 Statements
Limitations. The present study focuses on disentangling the color and texture elements in a training-free manner, and performing stylized image generation using these elements. This method can offer a flexible and customizable image generation for art creators and designers. However, when applying to content-consistency stylization, the consistency generation ability of this work can be further improved. In our future work, we will explore deeper with the consistency and expand our research into content-consistency stylization.
Broader Impacts. SADis enhances the flexible stylization capability in text-to-image synthesis by disentangling the color and texture elements. However, it also carries potential negative implications. It could be used to generate false or misleading images, thereby spreading misinformation. If SADis is applied to generate images of public figures, it poses a risk of infringing on personal privacy. Additionally, the automatically generated images may also touch upon copyright and intellectual property issues.
Ethical Statement. We acknowledge the potential ethical implications of deploying generative models, including issues related to privacy, data misuse, and the propagation of biases. All models used in this paper are publicly available. We will release the modified codes to reproduce the results of this paper. We also want to point out the potential role of customization approaches in the generation of fake news, and we encourage and support responsible usage.
Reproducibility Statement. To facilitate reproducibility, we will make the entire source code and scripts needed to replicate all results presented in this paper available after the peer review period. We will release the code for the novel color metric we have introduced. We conducted all experiments using publicly accessible datasets. Elaborate details of all experiments have been provided in the Appendices.
Appendix S2 Image-Prompt Additivity
Figure S1: By subtracting or adding the object image embeddings from the input image embeddings, we can effectively remove or add the object to the scene, although some degree of identity information for the person is also diminished.
Figure S2: (Left) By subtracting the grayscale image embeddings from the color image embeddings, we can effectively remove the texture information and isolate the color information. Combined with the prompt “A photo of a bird” we can generate images in the same color schemes as the input images. (Right) By subtracting the bluish image embeddings with the grayscale embeddings for random generated images with 100 prompts, we visualize the subtracted embedding via PCA decomposition. We observe them gathering around the pure blue image. The generation with these subtracted embeddings further prove the consistency.
Figure S3: By subtracting the color image embeddings with the grayscale embeddings for random generated images with 100 prompts, we visualize the subtracted embedding via PCA decomposition. We observe them gathering around the color reference image. The generation with these subtracted embeddings further prove the consistency.
S2.1 Broader Property of Image-Prompt Additivity
We demonstrate the broader property of Image-Prompt Additivity in Fig. S1. For instance, in Fig. S1-(a), subtracting the embedding of a hat from the embedding of a person wearing a hat results in the generation of a person without a hat. Similarly, in Fig. S1-(b), adding the embedding of glasses to a person results in the generation of that person wearing glasses. We refer to this phenomenon as Image-Prompt Additivity. We hypothesize this property originates from the image-text paired training in CLIP models [55]. The training process endows the image branch with the additivity property inherent to the text embedding space [51, 33, 6]. Although the person identities are altered after embedding additivity manipulations, this approach demonstrates a promising property for enabling our training-free color-texture disentanglement. Building on this property, we develop our method, SADis, to extract color and texture information from reference images and effectively apply these attributes in T2I generation.
In future work, we aim to address the limitations of Image-Prompt Additivity by enhancing identity consistency after additivity manipulations. This could broaden the impact and applicability of this property, advancing its utility in the field of image generation.
S2.2 Extra Analysis on Image-Prompt Additivity
As another illustration of image prompt additivity, we construct a set of blue, red, and green images. For example, the blue set is constructed by setting the red and green color channels to zero of a set of 100 images (similarly, for the green and red set). We isolate the color representation by performing a feature subtraction between the blue image and its grayscale equivalent. We plot their projected embeddings as blue dots in Fig. S2-(Right); note how these embeddings always maintain close to the pure blue embedding and are far from the green and red embeddings. Also, generated images with these subtraction blue color embeddings keep the bluish color palette as expected (see the car and t-shirt). An extended analysis is presented in Fig. S3, where we apply similar manipulations using colorful images as color references. The successful clustering further confirms that the resulting embeddings effectively preserve essential color information.
Appendix S3 Implementation details
We develop our method, SADis, based on the SDXL model [53], which is among the leading open-source T2I generative models available. To construct the experimental datasets, we randomly select 40 images from the WikiArt dataset [64] and StyleDrop [62] image collections. Each image can be used either as a color reference or a texture reference in the experiments. All input images are resized to $512\times 512$ before feeding into models. Also, for quantitative comparisons, 20 images are designated as the color reference set, while the remaining 20 serve as the texture reference set. Additionally, we sample 10 content prompts from StyleDrop, yielding a total of 4000 stylized images for each method for comparison. Since some comparison methods (such as SDXL [53] and Artist [39]) require text prompts as style controls for image generation, we employ the state-of-the-art vision-language model GPT-4o as the image captioning tool to generate precise color and texture prompts from the reference images. Subsequently, the content, color, and texture prompts are concatenated in the format: {content prompt}, {texture prompt} in the color {color prompt}. As for the comparison models (like DEADiff [54], IP-Adapter [75], Instantstyle [67], DreamStyler [3], StyleDrop [62], and CSGO [73]) that require text prompts for color control, the text prompt is constructed as: {content prompt}, in the color {color prompt}.
Figure S4: Some color and texture prompts generated by GPT-4o for comparison methods in the main paper for qualitative evaluation.
We further include the details of each evaluation metric. CLIP Score [55] is used to evaluate the semantic alignment between the text prompt and the generated image. We calucate the CLIP Score according to configuration of the T2I-CompBench repository [34]. For the color alignment evaluation, we use MS-SWD [25], color histogram distance (C-Hist), and GPT4o to calculate the color attribute similarity between the color reference and the generated image. MS-SWD [25]: since the generated imgae are usually not spatially aligned with the color reference image, we use MS-SWD to better evaluate the color attribute alignment according to the default setting of their repository [25]. C-Hist: we first compute the RGB histograms of the color reference and generated images. Afterwards, the Bhattacharyya distance is used to measure the differences between their color histograms. GPT-4o score [1]: to comprehensively evaluate the color alignment performance, we furture adopt the multimodal model GPT-4o to compute the color alignment score between the generated image and the color reference image. Specifically, the GPT-4o metric is computed according to Fig. S5. Firstly, following the previous work[17], we extract dominant colors by ColorThief. Afterwards, we feed the extracted color names to GPT-4o as the reference color from the color image, and ask GPT-4o to give 1-5 points according to the criteria shown in Fig. S5.
Figure S5: GPT-4o metric for color alignment evaluation.
Appendix S4 WCT transformations formulas
Given a latent $z_{t}^{{}^{\prime}}\in\mathcal{R}^{C\times H\times W}$ and a color reference latent $z_{t}^{c}\in\mathcal{R}^{C\times H\times W}$ at timestep $t$, we adopt Whitening and Coloring Transforms (WCT) [44, 10, 31] to transform $z_{t}^{{}^{\prime}}$ to match covariance matrix of color reference latent $z_{t}^{c}$. There are two step for WCT: Whitening transform and Coloring transform.
Whitening transform: the latent $z_{t}^{{}^{\prime}}$ is firstly centered by subtracting its mean vector $m$, and then an uncorrelated latent $\hat{z_{t}}^{{}^{\prime}}$ is obtained by:
$\hat{z_{t}}^{{}^{\prime}}$ $=ED^{-1/2}E^{\text{T}}z_{t}^{{}^{\prime}},$ (4)
where $D$ denotes a diagonal matrix of $z_{t}^{{}^{\prime}}z_{t}^{{}^{\prime}\text{T}}\in\mathcal{R}^{C\times C}$ and $E$ is the corresponding orthogonal matrix of eigenvectors which satisfy $z_{t}^{{}^{\prime}}z_{t}^{{}^{\prime}\text{T}}=EDE^{\text{T}}$.
Coloring transform: it’s the reverse process of Whitening transform [32, 44]. Beforehand, the color reference latent $z_{t}^{c}$ is centered by subtracting its mean vector $m_{c}$. Subsequently, we obtain the transformed latent $z_{t}^{wct}$ which satisfies the desired correlations $z_{t}^{wct}{z_{t}^{wct}}^{\text{T}}=z_{t}^{c}{z_{t}^{c}}^{\text{T}}=I$:
$z_{t}^{wct}$ $=E_{c}D_{c}^{-1/2}E_{c}^{\text{T}}\hat{z_{t}}^{{}^{\prime}},$ (5)
where $D_{c}$ is a diagonal matrix with the eigenvalues of the covariance matrix $z_{t}^{c}z_{t}^{c\text{T}}$ and $E_{c}$ is the corresponding orthogonal matrix of eigenvectors. Finally, we re-center the WCT transformed latent $z_{t}^{wct}$ by adding the mean vector $m_{c}$ of the color reference latent $z_{t}^{c}$:
$z_{t}^{wct}$ $=z_{t}^{wct}+m_{c}.$ (6)
Appendix S5 Additional Experimental Results
In Fig. S6, we include more color-texture disentanglement examples of our method SADis, which is under the Disentangled Stylized Image Generation (DisIG) scenario.
Figure S6: Additional experimental results of SADis.
Image-based stylization. Furthermore, our method, SADis, can be seamlessly integrated with ControlNet [77] to enable image-based stylized generation, as demonstrated in Fig. S7. Here, we adopt ControlNet (Canny) as the base pipeline, setting its conditioning scale to 0.6. All other hyper-parameters are kept consistent with those of SADis. This integration broadens the application scenarios of our proposed approach.
Figure S7: Our work is compatible with ControlNet to achieve image-based stylization.
Color and material transfer. Given color and material references from separate images, our method, SADis, is also capable of disentangling color and material elements, and provides precise controls over color and materials in T2I generation. The color and material transfer results are shown in Fig. S9. This capability of SADis highlights its potential applications in artistic creation and industrial design. We also desire to note that, the term “color palette” typically refers to the overall color scheme of an image, not specific local regions. Our SADis does not control the color palette of individual objects in general. However, Fig. S9 demonstrates controlled color and material transfer using a masking mechanism with our method SADis.
Figure S8: SADis can control color and texture generation separately using reference images.
Figure S9: SADis effectively disentangles color and material elements from separate images, enabling precise control over color and material in image generation.
Robust to color variations. As shown in Fig.S10 and Fig. S11, we vary the saturation and illuminant continuously. To be specific, for saturation, we developed a Python-based saturation adjustment tool utilizing HSV (Hue, Saturation, Value) color space transformation. The tool linearly modifies the saturation channel using factors of [0.2, 0.6, 1.0, 1.5, 3.0], respectively, before converting the images back to RGB format. The results of saturation variations are shown in Fig. S10. Regarding the color of the illuminant, we modify images by applying calibrated RGB channel multipliers. For warm temperatures (red) as shown in Fig. S11, red increases by 30%, green reduces by 10%, and blue by 20%. The intensity of these adjustments can be controlled (0-1 range), enabling fine-grained color temperature manipulation. We present 3 increasing intensities in Fig. S11. Our method SADis is able to generate images with smooth change along with the saturation and illuminant color variations, while the other method InstantStyle [67] fails. That further proves the robustness of our method SADis for Disentangled Stylized Image Generation (DisIG).
Figure S10: Compared to InstantStyle, SADis preserves the content more effectively under saturation variations.
Figure S11: Compared to InstantStyle, SADis preserves the content more effectively under illuminant variations.
Each component works independently. The T2I generations in Fig. S8-left provide evidence that the color and texture branches in SADis function independently. Additionally, the visualization for our ablatio study in  Fig. S8 further demonstrates improved color alignment with minimal texture degradation, achieving a good trade-off. However, the trade-off between texture and color alignment is optional for users, allowing them to adjust the balance based on the requirements of specific application scenarios.
Appendix S6 Ablation studies on RegWCT
RegWCT scales. Given a latent $z_{t}^{{}^{\prime}}$ at timestep $t$, the color rectified latent $z_{t}$ after RegWCT is presented as:
$z_{t}=\left(1-\omega\right)z_{t}^{{}^{\prime}}+\omega\cdot\mathbf{RegWCT}(z_{t}^{{}^{\prime}}).$ (7)
Here, we ablate the balancing factor $\omega$, as shown in Tab. S1. With the increasing scale $\omega$ of RegWCT, the color score of SADis improves significantly, while texture preservation remains largely unaffected. Based on the ablation results in Tab. S1, we set $\omega$ to 0.5 to achieve a balanced performance between color and texture alignment.
RegWCT Scale 0 0.3 0.5 0.7 1.0
Color MS-SWD (↓) 8.05 5.65 5.57 5.23 5.18
C-Hist (↓) 1.06 0.98 0.96 0.89 0.889
Texture KID (↓) 0.1077 0.1092 0.1076 0.1094 0.1081
CLIP-I (↑) 0.747 0.744 0.743 0.741 0.740
Table S1: Ablation studies of RegWCT scales. considering the trade-off of texture and color alignment, the RegWCT scale is set as 0.5 by default.
Timestep intervals. The ablation study on different timesteps to apply RegWCT is shown in Fig. S12-(a). During the denoising process, applying RegWCT to the latent enhances the color alignment performance but sacrifices texture preservation. As observed in Fig. S12-(a), the early stages of denoising contribute more to the color than the latter stages, which is also supported by previous works [17, 78, 66]. Therefore, we only apply RegWCT during the early stages of the denoising, specifically within $[0.8T,0.6T]$ to achieve a balance between the color and texture generation.
Figure S12: Ablation studies on applying RegWCT to different timestep intervals (left). Ablation studies on the scale $\lambda$ of noise injection (right) during applying RegWCT. By default, $\lambda$ is set to 0.01.
Figure S13: Results of different scales of color embedding $\mathbf{Emb}_{clr}$. Here, the scale of $\mathbf{Emb}_{tx}$ is fixed as 1.
$\gamma$ 0 0.001 0.003 0.005 0.007 0.009 0.011
Color MS-SWD (↓) 5.70 5.71 5.57 5.36 5.22 5.21 5.20
C-Hist (↓) 1.01 1.017 0.962 0.939 0.897 0.887 0.861
Texture KID (↓) 0.0986 0.1022 0.1076 0.1150 0.1215 0.1316 0.1309
CLIP-I (↑) 0.759 0.755 0.743 0.736 0.730 0.723 0.713
Table S2: Ablation studies of scaling factor $\gamma$. $\gamma$ is set to 0.003 according to the ablation results.
Scale $\lambda$ of noise injection. The ablation study on the noise scale $\lambda$ is illustrated in Fig. S12-(b). Without noise injection ($\lambda=0$), applying WCT improves color alignment but also causes great texture degradation for DisIG. This issue is alleviated by injecting the specific degree of latent noise $z_{T}$, which is demonstrated in Fig. S12-(b). To achieve a balanced color and texture alignment, we set $\lambda$ to 0.01 in the experiments based on the results shown in Fig. S12-(b).
Appendix S7 Ablation studies on CTE.
Scale $\gamma$ of SVD. The ablation results of the scaling factor $\gamma$ is presented in Tab. S2. As the scaling factor $\gamma$ increases, the color scores (such as MS-SWD and C-HIST) improve, with the sacrifice of the slightly texture performance degradation (revealed by CLIP-I and KID scores). Setting $\gamma=0.003$ achieves a balance trade-off between color and texture alignment.
Color-texture scales. As shown in Fig. S13, we fix the scale of the texture embedding $\mathbf{Emb}{tx}$ to 1 while varying the scale of the color embedding $\mathbf{Emb}{clr}$ from 0 to 1.8. With a greater weight for the color embedding, the color score improves. However, it is worth noting that when the scale of the color embedding is set too high, it may negatively impact texture preservation and introduce artifacts that are not derived from either the color or texture reference images.